OpenStudy (anonymous):

differentiate: a^4y=(x^2+a^2)^3/2

OpenStudy (lgbasallote):

with respect to what??

OpenStudy (anonymous):

y

OpenStudy (anonymous):

assuming a is constant. Let \[u=x^2+a^2\] \[y=\frac{u^{3}}{a^4}\] now find \[\frac{dy}{du}=\frac{3\sqrt{u}}{2a^4}\] and \[\frac{du}{dx}=2x\] now use chain rule \[\frac{dy}{dx}=\frac{dy}{du}*\frac{du}{dx}\]

OpenStudy (anonymous):

i don't understand .. huhu

OpenStudy (lgbasallote):

assuming a is not constant.... \[4a^3 da dy = \frac 32 (x^2 + a^2)^{1/2} (2x dx + 2ada)\] lol

OpenStudy (anonymous):

huh? i don't understand ..

OpenStudy (lgbasallote):

you werent supposed to lol. that was just a joke

OpenStudy (lgbasallote):

"lol" usually indicates "not serious"

OpenStudy (anonymous):

haay .. tsk ..